Circular No. 8247

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2003 UY_{275} (LINEAR) An apparently asteroidal object discovered by the LINEAR project

(discovery observation below from MPS 90712; other observations on MPS 91955, orbital elements on MPEC 2003-W49) has been found to show cometary appearance on CCD images taken with the Mt. Hopkins 1.2m reflector on Nov. 30.25 UT by C. W. Hergenrother; his co-added 1200-s R-band exposures show a highly condensed 16′′ coma and a narrow tail 100′′ long in p.a. 280^{o} (mag 18.5 determined by T. B. Spahr). Also, R. S. McMillan noted the object as diffuse in Spacewatch incidental observations made on Nov. 30.4.

2003 UT  α_{2000}  δ_{2000}  Mag.
Oct. 29.32233  4^{h}57^{m}33.^{s}53  +26^{o}23′38′′.1  20.2

Additional astrometry, the following improved orbital elements from observations Oct. 5–Dec. 1, and an ephemeris appear on MPEC 2003-X17.

T = 2003 July 2.0831 TT  ω = 119.^{o}3746
e = 0.508882  Ω = 245.7513 }2000.0
q = 1.831486 AU  i = 16.3301

a = 3.729221 AU n^{o} = 0.1368601 P = 7.202 years

COMET P/2001 RG_{100} = 1979 O1

Further to IAUC 8244, S. Nakano (Sumoto, Japan) reports that comet P/2001 RG_{100} (LINEAR) is identical with C/1979 O1 = 1979h (Kowal; cf. IAUC 3395 and 3397; MPC 4904 and 4998), suspected at the time as being of short period but lost immediately after discovery. Nakano’s improved orbital elements, given on Nakano Note 995, are very similar to those for the 2002 and 1992 epochs on MPEC 2003-W74, and he also gives elements for a 1981 epoch (T = 1981 Sept. 14).

COMET C/2001 Q4 (NEAT) Visual total-magnitude estimates: Aug. 26.56 UT, 12.0 (D. A. J. Sear-

gent, Cowra, N.S.W., 0.25-m reflector); Sept. 1.82, 12.2 (A. Pearce, Nedlands, W. Australia, 0.41-m reflector); 17.56, 11.6 (Seargent, 25×100 binoculars); Oct. 2.13, 12.0 (A. Amorim, Florianopolis, Brazil, 0.14-m reflector); Nov. 3.27, 11.1 (Amorim); 11.43, 10.4 (Seargent); 20.02, 10.6 (Amorim); 27.45, 10.2 (Seargent).

2003 December 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green